A method is used to prevent rusting in ships and underground pipelines involves connecting the iron to a block of a more active metal such as magnesium. This method is called cathodic protection. How many moles of magnesium are present in 1 billion (1 * 109) atoms of magnesium?


1
Expert's answer
2021-04-25T06:16:16-0400

Solution:

Magnesium - Mg

One mole of any substance contains 6.022×1023 atoms/molecules.

Hence,

Moles of Mg = (1.0×109 atoms of Mg) × (1 mol / 6.022×1023 atoms) = 1.66 ×10-15 moles of Mg


Answer: 1.66 ×10-15 moles of magnesium (Mg) are present in 1 billion atoms of magnesium (Mg).